OpenStudy (anonymous):

[6.03] What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57

terenzreignz (terenzreignz):

What methods of solving systems of linear equations have you been taught/ do you want to use on this one?

OpenStudy (anonymous):

elimination?

terenzreignz (terenzreignz):

That works, but if I may, this particular problems seems to be *begging* to be solved by substitution, are you fine with that? :)

terenzreignz (terenzreignz):

*problem, not problems, jeez

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Ok, look at the second equation: 4x + y = 57 Whenever you have a variable already alone (ie it has no number beside it) that usually makes life easier for you. In this case, try to isolate y (in other words, put it on one side of the equation, and all the rest on the other side) Go ahead :D

OpenStudy (anonymous):

y=0

terenzreignz (terenzreignz):

No. I meant like this: (I'm going to take you through this step by step, ok? ) 4x + y = 57 Now add (-4x) to both sides 4x + y - 4x = 57 - 4x 4x - 4x + y = 57 - 4x y = 57 - 4x Get me so far? :)

OpenStudy (anonymous):

-16+-4y=-228 and yes its fine the step by step

terenzreignz (terenzreignz):

If I may ask, where did you get that equation?

OpenStudy (anonymous):

from the 4x+y=57

terenzreignz (terenzreignz):

Oh, so you multiplied everything by -4? I see no reason to do that, perhaps you should try multiplying by 2, instead (by the way, this is in no way involved with substitution method, this is elimination method, but if this is what you're comfortable with, then go right ahead :) )

OpenStudy (anonymous):

2(4x+y=57)2 8x+2y=114 -8 -8 2y=106 2/106= 53

terenzreignz (terenzreignz):

Ok, I see you have trouble with elimination method, would you like me to explain it to you, step by step? (Don't worry, I'm more than willing to do so :D )

OpenStudy (anonymous):

yes that would be fine and thank you so much

terenzreignz (terenzreignz):

I'll start from the VERY basics, ok? For instance, if we have a = b and c = d it follows that a + c = b + d, right? if you don't get it, I can explain further, if you wish

OpenStudy (anonymous):

no i dont get it im sorry

terenzreignz (terenzreignz):

Ok, well, a = b, right? since they're equal, we can add c to both sides and still have it equal, like so: a + c = b + c but c = d, so in any expression that has c, we can replace it with d, so... a + c = b + d (we replaced the second c with d, since they're equal anyway) Get it now?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

ok, so we have established that if a = b and c = d then a + c = b + d right? Well, your systems of linear equations isn't THAT much different you could say your a, b, c, and d just got fancier :) Let me show you what I mean:

terenzreignz (terenzreignz):

7x - 2y = 21 4x + y = 57 your "a" is 7x - 2y your "b" is 21 your "c" is 4x + y your "d" is 57 Do you get me so far?

OpenStudy (anonymous):

okay i get that part

terenzreignz (terenzreignz):

so, if we apply the same rules: 7x - 2y = 21 4x + y = 57 we can say that 7x - 2y + 4x + y = 21 + 57 right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

We can simplify it to 11x - y = 78 But this does not really help us The trick to solving by elimination is properly *manipulating* the equations Ready for me to show you?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

ok, back to our original pair of equations 7x - 2y = 21 4x + y = 57 What we want to do is to alter them a bit, so that when we add them up like we just did, one of the variables will cancel out, leaving us with a linear equation in one variable. Any idea how to do this?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Ok, look at them again 7x - 2y = 21 4x + y = 57 Let's look at 7x for now What would cancel out 7x? -7x would... but there's no immediate way to turn 4x into -7x, is there? Perhaps the -2y is easier What would cancel out -2y?

OpenStudy (anonymous):

7 would cancel out to be just x and no there is no way 4 can turn to a -7x

terenzreignz (terenzreignz):

You're about right in your second statement, the first one is rather vague, but what would cancel out -2y?

OpenStudy (anonymous):

i need to turn this test in half hour and i need to check the other questions too

terenzreignz (terenzreignz):

Ok, I'll demonstrate it for you, but try to learn it ok?

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

7x - 2y = 21 4x + y = 57 We take the second equation, and multiply everything by 2 2(4x + y) = 2(57) 8x + 2y = 114 We replace the second equation with this one 7x - 2y = 21 8x + 2y = 114 Now we add up all those on the left side, and all those on the right 7x - 2y + 8x + 2y = 21 + 114 15x = 135 (the y got cancelled out!) x = 9 (divided both sides by 15) So now that we know that x = 9, then we substitute it for the x in the first equation 7x - 2y = 21 7(9) - 2y = 21 63 - 2y = 21 - 2y = 21 - 63 - 2y = -42 y = 21 (dividing both sides by -2) So x=9 and y=21 To be safe, let's try it with the second equation as well 4x + y = 57 4(9) + 21 = 57 36 + 21 = 57 57 = 57 (it works! ) Thus, we're sure that x = 9 and y = 21 END Hope that helps And remember to study the process I went through to get to the answer

OpenStudy (anonymous):

one more question and that would be all [6.02] Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

terenzreignz (terenzreignz):

4x + y = 11 x + 2y = 8 Take the first equation and add -4x to both sides 4x + y - 4x = 11 - 4x y = 11 - 4x We have obtained a value for y in terms of x. Now we substitute that value for the y in the second equation: x + 2y = 8 x + 2(11 - 4x) = 8 x + 22 - 8x = 8 -7x = 8 - 22 -7x = -14 (divide both sides by -7) x = 2 Now we substitute 2 for x in any of the two given equations, in the first: 4x + y = 11 4(2) + y = 11 8 + y = 11 y = 11 - 8 y = 3 We have x = 2 and y = 3. Let's verify with the second equation: x + 2y = 8 2 + 2(3) = 8 2 + 6 = 8 8 = 8 (it works!) Thus, we're sure that x = 2 and y = 3 END PS: Please go through the process involved, it's very important :)